OpenStudy (anonymous):

write the equation of the circle in standard form. find the center,radius and intercepts. x^2+y^2+12x+10y+61=16 im soo confused

OpenStudy (raden):

it similar with : (x+6)^2 + (y+5)^2 = 4^2 now, u can determine the centre and its radius that circle, just look general form circle equation : (x-a)^2 + (y-b)^2 = r^2, with a,b are the centre and r is radius